Grade 12th passMagical Mathematics[Interesting Approach]

Divide 32 into five parts in a.p. such that the product of the extremes is to the product of means as 7:15

Meenu Sharma
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Let 5 nos. in A.P. are
a-2d, a-d, a, a+d, a+2d
Sum = 5a = 32
a = 32/5
 
a-2d/a+2d = 7/15
15a- 30d = 7a +14d
8a = 44d
8*32/5 = 44 d
d = 64/55 
Hence now you can get all the numbers